# Utilizing Large Language Models to Optimize the Detection and Explainability of Phishing Websites

### v1

@@ Abstract.
In this paper, we introduce PhishLang, an open-source, lightweight Large Language Model (LLM) specifically designed for phishing website detection through contextual analysis of the website. Unlike traditional heuristic or machine learning models that rely on static features and struggle to adapt to new threats, and deep learning models that are computationally intensive, our model utilizes the advanced language processing capabilities of LLMs to learn granular features that are characteristic of phishing attacks. Furthermore, PhishLang operates with minimal data preprocessing and offers performance comparable to leading deep learning tools, while being significantly faster and less resource-intensive. Over a 3.5 months testing period, PhishLang successfully identified approximately 26K phishing URLs, many of which were undetected by popular antiphishing blocklists, thus demonstrating its potential to aid current detection measures. We also evaluate PhishLang against several realistic adversarial attacks, and develop six patches that make it very robust against such threats. Furthermore, we integrate PhishLang with GPT-3.5 Turbo to create explainable blocklisting - warnings that provide users with contextual information into different features that led to a website being marked as phishing. Finally, we have open-sourced the PhishLang framework, and developed a Chromium-based browser extension and URL scanner website, which implement explainable warnings for end-users.
@@ 1. Introduction
Phishing attacks have resulted in numerous data breaches and cases of credential theft, causing financial losses exceeding \$52M and impacting over 300K users during the last year alone (cis, 2021; Tom Huddleston Jr., [n. d.]). These scams are largely successful due to sophisticated social engineering techniques (Downs et al., 2007) that enable the creation of websites imitating legitimate organizations. To counter these threats, researchers across industry and academia have focused on developing countermeasures that leverage unique phishing signals, such as URL characteristics, network patterns, heuristic features in source code , and advanced machine learning methods (Nguyen et al., 2013; Sreedharan and Mohandas, 2016). A more recent focus has been on the visual aspects of phishing sites (Akhawe and Felt, 2013; Mohammad et al., 2015), with several methods now using image-based content, such as screenshots, for phishing detection (Abdelnabi et al., 2020; Lin et al., 2021). Although these models demonstrate high detection rates in controlled environments, they often do not fare well against complex evasion techniques (Yuan et al., 2023; Basit et al., 2021). Keeping these models up-to-date also poses significant challenges, as attackers continuously develop evasion tactics. Updating models to counter new threats often requires the creation of revised ground truth datasets and the identification of novel feature sets or strategies. Additionally, the inherent complexity and resource demands of the more recent deep learning models, especially those that depend on image content - limit their practicality for large-scale, real-time deployment (Saha Roy et al., 2023; Oest et al., 2020b). Furthermore, their reliance on opaque, black-box approaches complicates explainability (Do et al., 2022), limiting opportunities for iterative improvement.
To address these challenges, we have developed PhishLang, an efficient and transparent phishing detection framework that harnesses the capabilities of Large Language Models (LLMs) which are trained to comprehend the context of potential phishing websites. By evaluating these sites against known phishing patterns, our framework avoids reliance on static feature sets. Instead, it utilizes the advanced linguistic skills of LLMs to identify subtle patterns in the source code. Our model not only operates with significantly lower resource requirements—using up to three times less memory and seventeen times less storage than traditional deep learning-based phishing detection models—but also performs better than several ML-based models at detecting new attacks, especially those which contain evasive features. It is also highly resilient against realistic adversarial attacks. We also utilize our model in conjunction with GPT 3.5 Turbo (OpenAI, 2023) to build “Explainable Blocklisting” - an approach that provides users with contextual information on why a phishing website was blocked. This can not only help users understand and engage with the protection mechanism more effectively but also educate them on recognizing phishing attempts independently when protections are unavailable or fail. Finally, we open-source our framework, PhishLang, and also provide a Chromium-based browser extension that can proactively block threats from within the users’ web browser, as well as an online scanning website, which can be used to scan new URLs on demand. The structure of the paper is as follows: Section 3 elaborates on our methodology for parsing source code to isolate critical tags associated with phishing functionality (Section 3.1). This technique significantly reduces the input data size for the Large Language Models (LLMs), enhancing both training and inference efficiency. We employ this method to develop our training ground truth from a vast collection of both phishing and benign websites from the open-source PhishPedia (Lin et al., 2021) dataset (Section 3.2). Following this, we describe our training process for LLMs using a “sliding window” technique that divides the parsed code into smaller chunks, optimizing the contextual information available to the model (Section 3.3), a technique that has been used to maximize LLM training efficiency (Chen et al., 2023). To strike a balance between processing speed and detection accuracy, we evaluate various leading open-source language models to select the most effective one for our purpose (Section 3.4). To interpret the prediction capabilities of our model, we also perform a Local Interpretable Model-Agnostic Explanations (LIME) analysis (Ribeiro et al., 2016), identifying the features that our model focuses on (Section 7.3). Section 4 details the development and implementation of our real-time phishing detection system, PhishLang, which operates on a continuous stream of live URLs from Certstream to identify new phishing threats (Section 4.1). To identify how PhishLang compares to other prevalent countermeasures, we benchmark our model’s performance against traditional deep-learning-based phishing models (Section 4.2), as well as established anti-phishing tools (Section 4.3, also focusing on the detection of evasive phishing categories. We then assess the robustness of our model against adversarial phishing websites by testing sixteen realistic attacks that manipulate the website’s source code and develop patches as countermeasures.  5). In Section 7, we introduce a novel “Explainable Blocklist” mechanism that utilizes our model in conjunction with GPT 3.5T to generate phishing warnings that include contextual explanations that can clarify why a website was flagged, thereby enhancing user comprehension of the detection process. This approach can not only bolster the credibility of phishing alerts but also deepen users’ understanding of different phishing techniques. The primary contributions of our work are:
@@
    We utilized a Large Language model to build a lightweight open-source phishing detection framework - PhishLang, that can not only outperform several popular machine learning-based models, but also provide a better trade-off between inference speed, space complexity, and performance, making it more viable for real-world implementations.
@@
    We open-source PhishLang and make it available as a browser extension (both available at https://tinyurl.com/vcjer5my) and a URL scanning website (https://freephish.org).
@@ 2. Related Work
LLMs in Content Moderation: Open-source LLMs have previously shown promise as content moderation tools in diverse areas of social computing and security. For example, fine-tuned BERT models are used to identify cyberbullying instances (Behzadi et al., 2021) and to discern toxic triggers on Reddit platforms (Chong and Kwak, 2022). Demicir (Demırcı et al., 2022) combined LSTM and GPT-2 techniques to spot elusive malware. Several of these papers have demonstrated how LLMs can be trained on small amounts of data, with minimal data processing and without the need for defining specific features, which in turn outperform several traditional machine learning-based methods. However, utilizing LLMs to identify online scams and malicious websites has been very limited. Existing work (Koide et al., 2023; Heiding et al., 2023) have predominantly focused on using proprietary LLMs like ChatGPT (ope, 2023), Claude (cla, 2023), and Perplexity (goo, 2023). These models, while powerful, possess a proprietary architecture, thus preventing potential enhancements or alterations by the broader research community. Their reliance on a cost-per-use API model (OpenAI, 2023) further impedes scalability— a critical component in the fight against harmful content (Siad, 2023). Moreover, Derner at al.  (Derner and Batistič, 2023) has pointed out how attackers can adversarially exploit these commercial LLMs to evade classifications, thus significantly impacting their protective capabilities. A notable exception towards utilizing LLMs for scam detection is Su et al. (Su and Su, 2023), who have attempted to use BERT to detect malicious websites, though the analysis is limited only to the URL of the website. Thus, the current state accentuates the pressing need for an efficient open-source model tailored for phishing website detection, which is also resilient against adversarial attacks.
Explainable warnings: Security systems often use warning dialogs to convey crucial information, particularly when blocking malicious content or changes in system configurations that could affect usability (Raja et al., 2011). However, repetitive and generic warnings provide minimal insight into actual threats, leading to user desensitization and ignored alerts (Voronkov et al., 2017; Akhawe and Felt, 2013; Krol et al., 2012). This issue is prevalent in antiphishing systems integrated into web browsers or installed as plugins (saf, 2020; mca, 2022), which typically indicate a dangerous website without explaining why it was flagged. Despite evidence that providing contextual information about threats enhances user understanding and trust (Desolda et al., 2023; Nguyen et al., 2021), this approach is largely unadopted in commercial antiphishing tools. This study thus explores using LLMs to automatically generate detailed contextual warnings about phishing threats. Our approach highlights specific features of a website and explains their malicious nature. By combining features detected by our local LLM model with the advanced language synthesis of GPT-3.5T, our system generates high-quality warnings that inform users about malicious features, their locations, and harmful implications, aiming to enhance user trust and understanding in a scalable manner.
@@ 3. Method
Large Language Models excel at comprehending the semantics of natural language, enabling them to detect subtle nuances and patterns in text (Zhang et al., 2020; Sarti and Nissim, 2022). This capability is particularly beneficial for understanding complex textual data. The flexibility of LLMs extends to the realm of source code analysis as well. Researchers have effectively utilized models like BERT to analyze source code snippets, creating embeddings that capture the contextual relationships within the code (Kanade et al., 2020), which have also proven useful for applications such as code completion (Ciniselli et al., 2021). Building on this foundation, LLMs have been adapted for malware identification (Rahali and Akhloufi, 2023; Yesir and Soğukpinar, 2021). By training on a large repository of source code, these models can predict the characteristics of the syntax that should appear in a malicious file versus that in a benign file, detecting threats without the need for specific feature definitions or predefined rules during the training stage. We use a similar methodology for building our phishing detection model, which is capable of anticipating the necessary content (text or syntax) in a website to classify it as phishing, both at a macro level (the overall website context) and a micro level (specific code snippets). More specifically, if any part of a website’s source code - or a combination thereof—resembles that of known phishing sites, it is flagged as phishing. Conversely, if it reflects the characteristics of known benign sites, it is marked as safe.
We break down the various processes in our architecture below:
@@ 3.1. Parsing the source code
Figure 2. Distritbution of: Libraries/Frameworks found in: A) Phishing websites, B) Benign Websites, and JS Main function calls found in: C) Phishing websites.
@@
Thus, to identify tags that will be relevant for phishing detection, we at first manually evaluated a set of 500 phishing websites from our training dataset (See Section 3.3), co-relating the source code of each website with its rendered website (as well as it’s screenshot) to identify several tags which are used to include actionable phishing objects, i.e., elements in the source code that are either involved in user interactions, execute some malicious functionality or contain deceptive text. We dedicate the proceeding text to listing the chosen tags and why they are relevant: We analyze Headings (h1, h2, h3) as they are used to grab attention and may contain misleading language on phishing sites. Paragraphs (p) contain the main textual content and are often employed to convey deceptive messages. Links (a) are critical because they can redirect users to malicious sites or mimic legitimate URLs. Lists (ul, ol, li) on phishing sites might outline deceptive instructions. Form Tags (form) are used to collect sensitive data, making them a strong indicator of phishing. We also use title tags which appear on the browser tab and attackers might insert misleading titles that closely mimic legitimate sites. For example, a phishing site might use a title such as ”Secure Banking Login” to deceive users into thinking they are accessing their bank’s official website, which might give the (phishing website) credibility. We also include Footer tags as benign websites often contain more links or information in the footer compared to phishing sites (Yuan et al., 2024b), since the former is trying to provide resources to enhance user experience, whereas phishing attacks may provide little to no information in the footer. Script Tags (script) are utilized for various malicious purposes, including capturing keystrokes or loading phishing content. Input Tags (input) in phishing sites are designed to illicitly gather personal information. Button Tags (button) can be manipulated to execute harmful actions such as downloading malware. Iframe Tags (iframe) allow the stealthy embedding of malicious content. Meta Tags (meta) related to redirection can automatically redirect users to fraudulent sites. Anchor Tags with JavaScript (a with JS) can run malicious scripts upon clicking. Image Map Tags (map, area) are used in phishing attempts to create misleading graphics or hidden links. We use these tags to build a parser that can automatically retrieve content enclosed within these key HTML tags, and each selected element is processed to extract its textual content and specific attributes. The output from this parsing process is formatted in a structured text representation. Each element is represented by its tag type, followed by its content and relevant attributes. Figure 1 illustrates a phishing website and its subsequently generated parsed representation. Even tags that do not contain any elements are preserved with ¡EMPTY¿ tag notations. This is because phishing sites frequently utilize empty ¡div¿ or ¡a¿ tags to simulate legitimacy, misleading the users. All text content in the tags is converted to lowercase since LLMs are often sensitive to case differences, which can be interpreted as different tokens. Also, the number of unique tokens that the model needs to learn is significantly reduced. To give further credence to our tag selection, we queried GPT-4 with 3,000 true positive samples from our training dataset (Section 3.2), asking it to identify tags crucial for phishing detection in these websites. Prior literature has evaluated GPT-4 as highly proficient at detecting phishing websites (Chataut et al., 2024; Koide et al., 2023), making it a suitable choice for our analysis. The model confirmed all our selected tags and additionally suggested the applet and img tags. We chose not to include the applet tag, as it is outdated and deprecated in HTML5, which is predominantly used by modern phishing websites (Oest et al., 2020a). Only 5 of the websites provided to GPT-4 and 42 in the entire dataset contained an applet tag. Similarly, we excluded the img tag because it typically contains unique image filenames or externally sourced images Our framework, based on a language model, is not equipped to process images.
As mentioned, our chosen tags focus on actionable areas of the webpage that cannot be removed without affecting functionality, which attackers need for their malicious intent (e.g., capturing credentials). However, in Section 5, we explore how content within these tags can be manipulated by adversarial perturbations, preventing our detection model from identifying malicious websites. We subsequently suggest countermeasures (patches) through parser modification and adversarial training in Section 5 to prevent these attacks.
3.2. Training data

### v2

# PhishLang A Lightweight, Client-Side Phishing Detection Framework using MobileBERT for Real-Time, Explainable Threat Mitigation

@@ Abstract.
In this paper, we introduce PhishLang, an open-source, lightweight language model specifically designed for phishing website detection through contextual analysis of the website. Unlike traditional heuristic or machine learning models that rely on static features and struggle to adapt to new threats, and deep learning models that are computationally intensive, our model leverages MobileBERT, a fast and memory-efficient variant of the BERT architecture, to learn granular features characteristic of phishing attacks. PhishLang operates with minimal data preprocessing and offers performance comparable to leading deep learning anti-phishing tools, while being significantly faster and less resource-intensive. Over a 3.5-month testing period, PhishLang successfully identified 25,796 phishing URLs, many of which were undetected by popular antiphishing blocklists, thus demonstrating its potential to enhance current detection measures. Capitalizing on PhishLang’s resource efficiency, we release the first open-source fully client-side Chromium browser extension that provides inference locally without requiring to consult an online blocklist and can be run on low-end systems with no impact on inference times. Our implementation does not only outperform prevalent (server-side) phishing tools, but is significantly more effective than the limited commercial client-side measures available. Furthermore, we study how PhishLang can be integrated with GPT-3.5 Turbo to create ”Explainable blocklisting” — which upon detection of a website, provides users with detailed contextual information about the features that led to a website being marked as phishing.
@@ 1. Introduction
Phishing attacks have led to numerous data breaches and credential theft incidents, resulting in financial losses surpassing \$52 million and affecting over 300,000 users in the past year alone (cis, 2021; Tom Huddleston Jr., [n. d.]). These scams are particularly effective due to sophisticated social engineering techniques (Downs et al., 2007) that allow attackers to convincingly imitate legitimate websites. In response, researchers have focused on developing countermeasures that leverage distinctive phishing indicators, such as URL characteristics, network behavior, heuristic features in source code, and advanced machine learning techniques (Nguyen et al., 2013; Sreedharan and Mohandas, 2016). Recently, there has been a growing emphasis on the visual characteristics of phishing sites (Akhawe and Felt, 2013; Mohammad et al., 2015), leading to the use of image-based methods, such as screenshots, for phishing detection (Abdelnabi et al., 2020; Lin et al., 2021). Despite achieving high detection rates in controlled environments, these models often encounter challenges in real-world applications due to sophisticated evasion tactics used by attackers (Yuan et al., 2023; Basit et al., 2021).
Keeping these models updated presents significant challenges, as attackers continuously develop new evasion tactics. Updating models to counter these threats often necessitates creating revised ground truth datasets and identifying novel feature sets or strategies. Additionally, the inherent complexity and resource demands of recent deep learning models—especially those that rely on image content—limit their practicality for large-scale, real-time deployment (Saha Roy et al., 2023; Oest et al., 2020b). Furthermore, the reliance on opaque, black-box approaches complicates explainability (Do et al., 2022), hindering opportunities for iterative improvement.
To address these issues, we introduce PhishLang, an efficient and transparent phishing detection framework utilizing MobileBERT, a lightweight language model. PhishLang employs advanced linguistic capabilities to identify subtle phishing patterns in the website source code, thus moving beyond static feature sets. Our model operates with significantly lower resource requirements—using up to almost seven times less memory than the next resource efficient model (DistilBERT), while outperforming several ML-based models in detecting new attacks, particularly those with evasive features. Unlike most ML-based phishing detection approaches, we also evaluate PhishLang against various realistic adversarial attacks and implement patches to enhance the model’s robustness. Additionally, we integrate our model with GPT-3.5 Turbo (OpenAI, 2023) to create “Explainable Blocklisting,” a feature that provides users with contextual explanations on why a phishing website was blocked. This approach not only helps users understand and engage with the protection mechanisms more effectively but also educates them on recognizing phishing attempts independently when such protections are unavailable or fail.
To the best of our knowledge, we have also developed the first open-source, Chromium-based browser extension that operates entirely on the client side, running locally on the user’s device without relying on an online blocklist, thus enhancing user privacy and eliminating blocklist-related latency. The PhishLang extension is designed to be highly efficient, with minimal memory usage, allowing it to run smoothly even on low-end hardware while delivering near-instant URL predictions. Despite being underexplored, client-side anti-phishing detection offers crucial benefits in phishing detection: it enhances user data privacy by keeping browsing data local, thus minimizing exposure risks, and eliminates the reliance on often delayed online blocklist updates (Oest et al., 2020a). Furthermore, client-side detection reduces network latency, providing faster threat responses, ensuring continued protection during server outages. Our implementation significantly surpasses existing client-side phishing solutions, such as Google Safe Browsing, which detects only 14% of phishing websites without server-side support (Pourmohamad et al., 2024).
The structure of the paper is as follows: Section 3 elaborates on our methodology for parsing source code to isolate critical tags associated with phishing functionality (Section 3.1). This technique significantly reduces the input data size for the language models, thereby improving both training and inference efficiency. We use this method to create our training ground truth from the open-source PhishPedia (Lin et al., 2021) dataset (Section 3.2.1). To balance processing speed with detection accuracy, we evaluate various leading open-source language models to identify the most effective one for our needs (Section 3.2.2).
Section 4 outlines the development and implementation of our real-time phishing detection system, PhishLang, which continuously monitors live URLs from Certstream to detect new phishing threats (Section 4.1). To benchmark our model’s effectiveness against existing measures, we also compare PhishLang’s performance with traditional deep learning-based phishing models (in Appendix), as well as (Section 9.1) established anti-phishing tools, with a focus on detecting evasive phishing tactics (Section 4.2). We then evaluate the robustness of our model against adversarial phishing websites by testing sixteen realistic attack scenarios that manipulate website source code, developing patches as countermeasures (Section 5). In Section 6, we introduce a novel “Explainable Blocklist” mechanism that pairs PhishLang with GPT-3.5 Turbo to generate phishing alerts that include contextual explanations, which can help users understand why a website was flagged. This approach not only enhances the credibility of phishing warnings but can also deepen users’ understanding of various phishing techniques. Finally, in Section 7, we highlight our client-side phishing detection extension. The primary contributions of our work are:
@@
    We utilized MobileBERT, a resource-efficient language model to build a lightweight open-source phishing detection framework - PhishLang, that can not only outperform several popular machine learning-based models, but also provide a better trade-off between inference speed, space complexity, and performance, making it more viable for real-world implementations.
@@
    We open-source our framework (https://github.com/UTA-SPRLab/phishlang) and implement PhishLang as a fully client-side Chromium browser extension with very little memory impact and can run on low-end hardware, that provides protection without online blocklists.
@@ 2. Related Work
Language models in Content Moderation: Open-source language models have previously shown promise as content moderation tools in diverse areas of social computing and security. For example, fine-tuned BERT models are used to identify cyberbullying instances (Behzadi et al., 2021) and to discern toxic triggers on Reddit platforms (Chong and Kwak, 2022). Demicir (Demırcı et al., 2022) combined LSTM and GPT-2 techniques to spot elusive malware. Several of these papers have demonstrated how LLMs can be trained on small amounts of data, with minimal data processing and without the need for defining specific features, which in turn outperform several traditional machine learning-based methods. However, utilizing these models to identify online scams and malicious websites has been very limited. Existing work (Koide et al., 2023; Heiding et al., 2023) have predominantly focused on using proprietary LLMs like ChatGPT (ope, 2023), Claude (cla, 2023), and Perplexity (goo, 2023). These models, while powerful, possess a proprietary architecture, thus preventing potential enhancements or alterations by the broader research community. Their reliance on a cost-per-use API model (OpenAI, 2023) further impedes scalability— a critical component in the fight against harmful content (Siad, 2023). Moreover, Derner at al.  (Derner and Batistič, 2023) has pointed out how attackers can adversarially exploit these commercial LLMs to evade classifications, thus significantly impacting their protective capabilities. A notable exception towards utilizing language models for scam detection is Su et al. (Su and Su, 2023), who have attempted to use BERT to detect malicious websites, though the analysis is limited only to the URL of the website. Thus, the current state accentuates the need for an efficient open-source model tailored for phishing website detection, which is also resilient against adversarial attacks.
@@ 3. Method
Large Language Models excel at comprehending the semantics of natural language, enabling them to detect subtle nuances and patterns in text (Zhang et al., 2020; Sarti and Nissim, 2022). This capability is particularly beneficial for understanding complex textual data. The flexibility of LLMs extends to the realm of source code analysis as well. Researchers have effectively utilized models like BERT to analyze source code snippets, creating embeddings that capture the contextual relationships within the code (Kanade et al., 2020), which have also proven useful for applications such as code completion (Ciniselli et al., 2021). Building on this foundation, LLMs have been adapted for malware identification (Rahali and Akhloufi, 2023; Yesir and Soğukpinar, 2021). By training on a large repository of source code, these models can predict the characteristics of the syntax that should appear in a malicious file versus that in a benign file, detecting threats without the need for specific feature definitions or predefined rules during the training stage. We use a similar methodology for building our phishing detection model, which is capable of anticipating the necessary content (text or syntax) in a website to classify it as phishing, both at a macro level (the overall website context) and a micro level (specific code snippets). More specifically, if any part of a website’s source code - or a combination thereof—resembles that of known phishing sites, it is flagged as phishing. Conversely, if it reflects the characteristics of known benign sites, it is marked as safe. We break down the processes in our architecture below:
@@ 3.1. Parsing the source code
Thus, to identify tags that will be relevant for phishing detection, we at first manually evaluated a set of 500 phishing websites from our training dataset (See Section 3.2), co-relating the source code of each website with its rendered website (as well as it’s screenshot) to identify several tags which are used to include actionable phishing objects, i.e., elements in the source code that are either involved in user interactions, execute some malicious functionality or contain deceptive text, that cannot be removed without affecting functionality needed by attackers need for their malicious intent (e.g., capturing credentials). We dedicate the proceeding text to listing the chosen tags and why they are relevant: We analyze Headings (h1, h2, h3) as they are used to grab attention and may contain misleading language on phishing sites. Paragraphs (p) contain the main textual content and are often employed to convey deceptive messages. Links (a) are critical because they can redirect users to malicious sites or mimic legitimate URLs. Lists (ul, ol, li) on phishing sites might outline deceptive instructions. Form Tags (form) are used to collect sensitive data, making them a strong indicator of phishing. We also use title tags which appear on the browser tab and attackers might insert misleading titles that closely mimic legitimate sites. For example, a phishing site might use a title such as ”Secure Banking Login” to deceive users into thinking they are accessing their bank’s official website, which might give the (phishing website) credibility. We also include Footer tags as benign websites often contain more links or information in the footer compared to phishing sites (Yuan et al., 2024b), since the former is trying to provide resources to enhance user experience, whereas phishing attacks may provide little to no information in the footer. Script Tags (script) are utilized for various malicious purposes, including capturing keystrokes or loading phishing content. Input Tags (input) in phishing sites are designed to illicitly gather personal information. Button Tags (button) can be manipulated to execute harmful actions such as downloading malware. Iframe Tags (iframe) allow the stealthy embedding of malicious content. Meta Tags (meta) related to redirection can automatically redirect users to fraudulent sites. Anchor Tags with JavaScript (a with JS) can run malicious scripts upon clicking. Image Map Tags (map, area) are used in phishing attempts to create misleading graphics or hidden links. We use these tags to build a parser that can automatically retrieve content enclosed within these key HTML tags, and each selected element is processed to extract its textual content and specific attributes. The output from this parsing process is formatted in a structured text representation. Each element is represented by its tag type, followed by its content and relevant attributes. Figure 1 illustrates a phishing website and its subsequently generated parsed representation. Even tags that do not contain any elements are preserved with ¡EMPTY¿ tag notations. This is because phishing sites frequently utilize empty ¡div¿ or ¡a¿ tags to simulate legitimacy, misleading the users. All text content in the tags is converted to lowercase since LLMs are often sensitive to case differences, which can be interpreted as different tokens. Also, the number of unique tokens that the model needs to learn is significantly reduced. To give further credence to our tag selection, we queried GPT-4 with 3,000 true positive samples from our training dataset (Section 3.2.1), asking it to identify tags crucial for phishing detection in these websites. Prior literature has evaluated GPT-4 as highly proficient at detecting phishing websites (Chataut et al., 2024; Koide et al., 2023), making it a suitable choice for our analysis. The model confirmed all our selected tags and additionally suggested the applet and img tags. We chose not to include the applet tag, as it is outdated and deprecated in HTML5, which is predominantly used by modern phishing websites (Oest et al., 2020a). Only 5 of the websites provided to GPT-4 and 42 in the entire dataset contained an applet tag. Similarly, we excluded the img tag because it typically contains unique image filenames or externally sourced images Our framework, based on a language model, is not equipped to process images.
3.2. Training and Inference
3.2.1. Groundtruth